# Gravity as a topological gauge theory

### v3

@@ 3.2 Characteristic poylynomials for $G=SO(4,2)/\{\pm I\}$ and associated action
$\begin{cases}\alpha_{1}=\dfrac{4k}{\ell\ell^{\prime}}\alpha_{4}=\dfrac{2\ell\ell^{\prime}}{k}\alpha_{6}\\ \alpha_{2}=\dfrac{k(r+2y)}{4\pi^{2}\ell\ell^{\prime}}=\dfrac{2k(r+2y)}{(r+y)\ell\ell^{\prime}}\alpha_{3}=-(1+\dfrac{2y}{r})\alpha_{5}\\ \alpha_{1}=\dfrac{e}{r+2y}\alpha_{2}\end{cases}$ (5)
@@
$S_{G}[\varpi]=\int_{\mathscr{M}}$ $\Bigl{(}\overbrace{\dfrac{k}{4\pi^{2}\ell\ell^{\prime}}\bigl{(}\dfrac{e}{2}R^{ab}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}+(r+2y)R^{ab}\wedge\beta_{a}\wedge\beta_{b}\bigr{)}}^{Holst}+\overbrace{\dfrac{2ek^{2}}{(4\pi\ell\ell^{\prime})^{2}}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}}^{Cosmological~{}constant}$
$\qquad+\,\overbrace{\dfrac{r+y}{8\pi^{2}}R^{ab}\wedge R_{ab}}^{Pontrjagin}+\overbrace{\dfrac{e}{2(4\pi)^{2}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}\overbrace{-\dfrac{rk}{4\pi^{2}\ell\ell^{\prime}}(T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{a}\wedge\beta_{b})}^{Nieh-Yan}$
$\qquad\overbrace{-\,\dfrac{rk}{2\pi^{2}\ell\ell^{\prime}}\eta_{ab}T^{a}\wedge\lambda\wedge\beta^{b}}^{Torsion-dilation~{}interaction}\overbrace{-\dfrac{r}{4\pi^{2}}d\lambda\wedge d\lambda}^{Dilation~{}kinetic~{}term}\Bigr{)}.$ (6)
Choosing the symmetric decomposition $\mathfrak{g}=\mathfrak{g}_{0}\oplus\mathfrak{m}$ with $\mathfrak{m}=\mathfrak{g}_{1}\oplus\mathfrak{g}_{-1}$ and $\mathfrak{g}_{0}=\mathfrak{co}(3,2)$ such that the Killing metric is computed to be:
$K_{\mathfrak{g}}(\varphi,\varphi^{\prime})=\overbrace{-4\eta_{ac}\eta_{bd}\varphi_{\mathfrak{so}}^{ab}\varphi_{\mathfrak{so}}^{{}^{\prime}cd}+8\varphi_{\mathbb{R}}\varphi^{\prime}_{\mathbb{R}}}^{K_{0}}+\overbrace{8\eta_{ab}(\varphi_{-1}^{a}\varphi_{1}^{{}^{\prime}b}+\varphi_{1}^{a}\varphi_{-1}^{{}^{\prime}b})}^{K_{\mathfrak{m}}}$
and after fixing $\alpha=\dfrac{k}{2}\beta$, one gets for the metrics:
$h(\gamma,\gamma^{\prime})=\kappa K_{\mathfrak{g}}(\gamma,\gamma^{\prime})$ (7)
$g(X,Y)=\zeta(\varpi_{\mathfrak{m}}^{*}h)(X,Y)=\kappa\zeta(\varpi_{\mathfrak{m}}^{*}K_{\mathfrak{m}})(X,Y)=\dfrac{8k\kappa\zeta}{\ell\ell^{\prime}}\,\eta\bigl{(}\beta(X),\beta(Y)\bigr{)}.$ (8)
If one considers the usual covariant derivative $D_{\mathfrak{so}}=(\partial_{\mu}+\dfrac{1}{4}A_{ab,\mu}\gamma^{a}\gamma^{b})dx^{\mu}$, the matter action similar to (3) is then given by  [^9] Where we set $\kappa^{2}\zeta=\tfrac{(\ell\ell^{\prime})^{2}}{192(2\ell^{\prime}+k\ell)k}$ such that the overall factor is arranged to stick with the Dirac action given in [7, p.197].:
$S_{M}[A,\beta]$ $=\int_{\mathscr{M}}h^{\varepsilon}\bigl{(}i\psi,h_{ab}\gamma^{a}\varpi_{\mathfrak{m}}^{b}\wedge*D_{\mathfrak{so}}\psi\bigl{)}=\mathrm{Re}S_{D}\,.$ (9)
@@
Thus, a more natural matter action $S^{\prime}_{M}$ with covariant derivative $D_{0}=d+\Phi_{\varepsilon}(\varpi_{0})=d+\Phi_{\varepsilon}(A-\lambda\mathds{1}_{4})=(\partial_{\mu}+\dfrac{1}{4}A_{ab,\mu}\gamma^{a}\gamma^{b}-\dfrac{1}{2}\lambda_{\mu})dx^{\mu}$ taking into account the full action of the $\mathfrak{g}_{0}=\mathfrak{co}(3,2)$-valued part of the connection $\varpi_{0}$ is:
$S^{\prime}_{M}[A,\beta,\lambda]$ $=\int_{\mathscr{M}}h^{\varepsilon}\bigl{(}i\psi,h_{ab}\gamma^{a}\varpi_{\mathfrak{m}}^{b}\wedge*D_{0}\psi\bigl{)}$ (14)
$=\overbrace{S_{M}}^{\mathrm{Re}S_{D}}-\dfrac{1}{2}\int_{\mathscr{M}}h^{\varepsilon}\bigl{(}i\psi,h_{ab}\gamma^{a}\varpi_{\mathfrak{m}}^{b}\wedge*(\lambda\psi)\bigl{)}.$ (15)
The study of the full action $S^{\prime}_{T}=S_{G}+S^{\prime}_{M}$ is left for future work. In the following we will restrict ourselves to the study of the action $S_{T}=S_{G}+S_{M}$.
@@
$\dfrac{\delta\mathscr{L}_{G}[A,\beta,\lambda]}{\delta\beta^{c}}$ $=\dfrac{ek}{4\pi^{2}\ell\ell^{\prime}}(R^{ab}+\dfrac{2k}{\ell\ell^{\prime}}\beta^{a}\wedge\beta^{b})\wedge\beta^{d}\varepsilon_{abcd}+\dfrac{k}{2\pi^{2}\ell\ell^{\prime}}(2(r+y)R^{ab}\wedge\beta_{b}\delta_{ac}+rT_{c}\wedge\lambda-r\beta_{c}\wedge d\lambda)$
@@
$\dfrac{\delta\mathscr{L}_{G}[A,\beta,\lambda]}{\delta A^{ab}}$ $=\dfrac{ek}{4\pi^{2}\ell\ell^{\prime}}T^{c}\wedge\beta^{d}\varepsilon_{abcd}+\dfrac{k}{\pi^{2}\ell\ell^{\prime}}\bigl{(}(\tfrac{r}{2}+y)T_{a}+\dfrac{r}{4}\lambda\wedge\beta_{a}\bigr{)}\wedge\beta_{b}=\dfrac{1}{2}\mathfrak{s}_{ab}=-\dfrac{\delta\mathscr{L}_{M}[A,\beta,\lambda]}{\delta A^{ab}}$ (17)
$\dfrac{\delta\mathscr{L}_{G}[A,\beta,\lambda]}{\delta\lambda}$ $=-\dfrac{rk}{2\pi^{2}\ell\ell^{\prime}}T^{a}\wedge\beta_{a}=0=-\dfrac{\delta\mathscr{L}_{M}[A,\beta,\lambda]}{\delta\lambda}\ .$ (18)
Equation (18) implies either $r=0$ or $T^{a}\wedge\beta_{a}=0$.
For $r=0$, one then recovers the usual equations of motion associated to Einstein-Cartan gravity with a Holst action and bare cosmological constant term, and torsion $T$ linked to the spin density of matter.
In the other case, using the Hodge star operator lets us show that $T^{a}\wedge\beta_{a}=0$ is equivalent to requiring the axial vector part of torsion to vanish ($\tilde{t}^{a}=\dfrac{1}{6}T^{b}_{~{}\mu\nu}\varepsilon^{a~{}\mu\nu}_{~{}b}=0$) as is described in [15, Table 3]. The end result being that the new torsion $T^{a}_{~{}bc}=Z^{a}_{~{}bc}+\delta^{a}_{[b}t^{\vphantom{a}}_{c]}$ can be expressed solely in terms of $20$ independent components instead of the usual $24$. Here $Z^{a}_{~{}bc}$ is a $3$-indices tensor and $t_{c}$ is a $4$-dimensional vector with respectively $16$ and $4$ independent components.
@@
$\dfrac{ek}{4\pi^{2}\ell\ell^{\prime}}(R^{ab}+\dfrac{2k}{\ell\ell^{\prime}}\beta^{a}\wedge\beta^{b})\wedge\beta^{d}\varepsilon_{abcd}+\dfrac{k}{2\pi^{2}\ell\ell^{\prime}}\bigl{(}2(r+y)R^{ab}\wedge\beta_{b}\delta_{ac}-r\beta_{c}\wedge d\lambda\bigr{)}$ $=\tau_{c}$ (19)
$\dfrac{rk}{4\pi^{2}\ell\ell^{\prime}}\lambda\wedge\beta_{a}\wedge\beta_{b}$ $=\dfrac{1}{2}\mathfrak{s}_{ab}$ (20)
@@
where equation (20) links spin density to the scalar field:
$\lambda_{c}=\lambda_{\mu}\beta^{-1\,\mu}_{~{}~{}\,\,\;\;\;c}=\dfrac{\pi^{2}\ell\ell^{\prime}}{3rk}\mathfrak{s}^{ab}_{~{}~{}abc}\ .$ (22)
Using the relation $\gamma=\dfrac{2\alpha_{1}}{2\alpha_{2}-\alpha_{5}}$, given in [8, p13], we identify $\gamma=\dfrac{2e}{3r+4y}$, with $\gamma$ the Barbero-Immirzi parameter.
Provided we set $e=\dfrac{4\pi^{2}\alpha_{1}\ell\ell^{\prime}}{k}=\dfrac{3\pi}{2\Lambda_{0}G}$ and $\ell\ell^{\prime}=-\dfrac{6k}{\Lambda_{0}}$ where $\alpha_{1}=-\dfrac{1}{16\pi G}$, applying the Hodge star * to (19) and replacing $\lambda$ according to (22) gives us the Einstein equations modified by the Barbero-Immirzi part of the action with bare cosmological constant $\Lambda_{0}$ and with matter sources given by the energy-momentum tensor and a particular combination of derivatives of the spin density of matter (the latter resulting from the presence of the dilation 1-form $\lambda$ of the connection):
$G_{kc}+(\dfrac{\Lambda_{0}Gr}{6\pi}+\dfrac{1}{2\gamma})R^{ab}_{~{}~{}rs}\varepsilon^{rs}_{~{}~{}bk}\delta_{ac}+\Lambda_{0}\eta_{kc}=-\dfrac{1}{2\alpha_{1}}(\tau_{ck}-\dfrac{1}{12}\partial_{\mu}\mathfrak{s}^{ab}_{~{}~{}ab\nu}\varepsilon^{\mu\nu}_{~{}~{}\;ck}).$ (23)
@@
$G_{kc}+(\dfrac{\Lambda_{0}Gr}{6\pi}+\dfrac{1}{2\gamma})R^{ab}_{~{}~{}rs}\varepsilon^{rs}_{~{}~{}bk}\delta_{ac}+\Lambda\eta_{kc}=-\dfrac{1}{2\alpha_{1}}(\tau_{M,ck}-\dfrac{1}{12}\partial_{\mu}\mathfrak{s}^{ab}_{~{}~{}ab\nu}\varepsilon^{\mu\nu}_{~{}~{}\;ck})$ (24)
with $\Lambda=\Lambda_{0}+\Lambda_{\text{vac}}=\overbrace{-\dfrac{6k}{\ell\ell^{\prime}}}^{\Lambda_{0}}\overbrace{-\dfrac{\rho_{\text{vac}}}{2\alpha_{1}}}^{\Lambda_{\text{vac}}}=-\dfrac{6k}{\ell\ell^{\prime}}-\dfrac{2\pi^{2}\ell\ell^{\prime}\rho_{\text{vac}}}{ek}=\Lambda_{0}+8\pi G\rho_{\text{vac}}$ the effective cosmological constant, and $\Lambda_{\text{vac}}$ the vacuum contribution to the cosmological constant.
@@
$G_{kc}$ $+(\dfrac{\Lambda_{0}Gr}{6\pi}+\dfrac{1}{2\gamma})R^{ab}_{~{}~{}rs}\varepsilon^{rs}_{~{}~{}bk}\delta_{ac}+\Lambda_{0}\eta_{kc}$
$=-\dfrac{1}{2\alpha_{1}}\Bigl{(}\tau_{ck}-\dfrac{1}{24}(2\eta_{cs}\beta^{s}_{\nu}\partial_{\mu}+T_{c,\mu\nu})\bigl{(}\mathfrak{s}^{ab}_{~{}~{}abr}+\dfrac{4(r+2y)}{\pi^{2}}T^{m}_{~{}rm}\bigr{)}\varepsilon^{\mu\nu r}_{~{}~{}~{}\,k}\Bigr{)}$ (25)
$\lambda_{c}$ $=\dfrac{1}{r}\bigl{(}\dfrac{\pi^{2}k}{3\ell\ell^{\prime}}\mathfrak{s}^{ab}_{~{}~{}abc}-8(r+2y)T^{a}_{~{}ca}\bigr{)}$ (26)

### v5

@@ 3.2 Characteristic poylynomials for $G=SO(4,2)/\{\pm I\}$ and associated action
$\begin{cases}\alpha_{1}=\dfrac{2k}{\ell\ell^{\prime}}\alpha_{4}=\dfrac{4\ell\ell^{\prime}}{k}\alpha_{6}\\ \alpha_{2}=\dfrac{ky}{4\pi^{2}\ell\ell^{\prime}}=\dfrac{2ky}{(r+y)\ell\ell^{\prime}}\alpha_{3}=-\dfrac{y}{r}\alpha_{5}\\ \alpha_{1}=\dfrac{e}{2y}\alpha_{2}.\end{cases}$ (8)
@@
$S_{G}[\varpi]=\int_{\mathscr{M}}$ $\Bigl{(}\overbrace{\dfrac{k}{4\pi^{2}\ell\ell^{\prime}}\big{(}\dfrac{e}{4}\underbrace{R^{ab}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}}_{Palatini}+yR^{ab}\wedge\beta_{a}\wedge\beta_{b}\big{)}}^{Holst}+\overbrace{\dfrac{ek^{2}}{32\pi^{2}(\ell\ell^{\prime})^{2}}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}}^{Bare~{}Cosmological~{}constant}$
$\qquad+\overbrace{\dfrac{r+y}{8\pi^{2}}R^{ab}\wedge R_{ab}}^{Pontrjagin}+\overbrace{\dfrac{e}{2(4\pi)^{2}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}\overbrace{-\dfrac{rk}{4\pi^{2}\ell\ell^{\prime}}(T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{a}\wedge\beta_{b})}^{Nieh-Yan}\Bigr{)}$
$\qquad\overbrace{-\dfrac{r+y}{4\pi^{2}}d\lambda\wedge d\lambda}^{Dilation~{}kinetic~{}term}\Bigr{)}.$ (9)
This action corresponds to (8) with the substitution $k/\ell^{2}\to k/(\ell\ell^{\prime})$ plus an additional kinetic term for the dilations. We will show in the next two subsections that in the presence of matter:
  • Action (3.2) gives back the equations of motion (9)-(10).
  • Adding the term $-\,\dfrac{vk}{2\pi^{2}\ell\ell^{\prime}}\eta_{ab}T^{a}\wedge\lambda\wedge\beta^{b}$ in the action (3.2) leads to a second source of curvature expressed in terms of the spin density of matter, torsion and their variations for the Einstein equations.
3.2.1 Fermionic matter field action
Choosing the symmetric decomposition $\mathfrak{g}=\mathfrak{g}_{0}\oplus\mathfrak{m}$ with $\mathfrak{m}=\mathfrak{g}_{1}\oplus\mathfrak{g}_{-1}$ and $\mathfrak{g}_{0}=\mathfrak{h}=\mathfrak{co}(3,1)$ we compute the Killing metric:
$K_{\mathfrak{g}}(\varphi,\varphi^{\prime})=\overbrace{-4\eta_{ac}\eta_{bd}\varphi_{\mathfrak{so}}^{ab}\varphi^{\prime}{}_{\!\!\mathfrak{so}}^{\,cd}+8\varphi_{\mathbb{R}}\varphi^{\prime}_{\mathbb{R}}}^{K_{0}}+\overbrace{\dfrac{8}{\ell\ell^{\prime}}\eta_{ab}(\varphi_{-1}^{a}\varphi^{\prime}{}_{\!\!1}^{\,b}+\varphi_{1}^{a}\varphi^{\prime}{}_{\!\!-1}^{\,b})}^{K_{\mathfrak{m}}}$ (10)
On the 4-dimensional submanifold $\mathscr{M}$ defined by the gauge fixing $\alpha=\dfrac{k}{2}\beta$, one has $\varpi_{\mathfrak{m}}=\beta^{a}M_{a}$ and one gets for the metrics
$\text{on}~{}\mathfrak{g}:$ $h(\varphi,\varphi^{\prime})=\kappa K_{\mathfrak{g}}(\varphi,\varphi^{\prime})$ (11)
$\text{on}~{}\mathscr{M}:$ $g(X,Y)=\zeta(\varpi_{\mathfrak{m}}^{*}h)(X,Y)=\kappa\zeta(\varpi_{\mathfrak{m}}^{*}K_{\mathfrak{m}})(X,Y)=\dfrac{8k\kappa\zeta}{\ell\ell^{\prime}}\,\eta\bigl{(}\beta(X),\beta(Y)\bigr{)}.$ (12)
If one considers the usual covariant derivative $D_{\mathfrak{so}}=(\partial_{\mu}+\dfrac{1}{4}A_{ab,\mu}\gamma^{a}\gamma^{b})dx^{\mu}$, the matter action similar to (3) is then given by  [^20] Where we set $\kappa^{2}\zeta=\tfrac{(\ell\ell^{\prime})^{2}}{384k^{2}}$ such that the overall factor is arranged to stick with the Dirac action given in [17, p.197].:
$\hskip-22.76219ptS_{M}[A,\beta]$ $=\int_{\mathscr{M}}h^{\varepsilon}\bigl{(}i\psi,h_{ab}\gamma^{a}\varpi_{\mathfrak{m}}^{b}\wedge*D_{\mathfrak{so}}\psi\bigl{)}=\int_{\mathscr{M}}h^{\varepsilon}\bigl{(}i\psi,\eta_{ab}\dfrac{64k^{2}\kappa^{2}\zeta}{(\ell\ell^{\prime})^{2}}\gamma^{a}\beta^{b}\wedge\tilde{*}D_{\mathfrak{so}}\psi\bigl{)}=\mathrm{Re}S_{D}\,.$ (13)
@@ 3.2.1 Fermionic matter field action
Thus, a more natural matter action $\tilde{S}_{M}$ with covariant derivative $D_{0}=d+\Phi_{\varepsilon}(\varpi_{0})=d+\Phi_{\varepsilon}(A-\lambda\mathds{1}_{4})=(\partial_{\mu}+\dfrac{1}{4}A_{ab,\mu}\gamma^{a}\gamma^{b}-\dfrac{1}{2}\lambda_{\mu})dx^{\mu}$ taking into account the full action of the $\mathfrak{g}_{0}=\mathfrak{co}(3,2)$-valued part of the connection $\varpi_{0}$ is:
$\tilde{S}_{M}[A,\beta,\lambda]$ $=\int_{\mathscr{M}}h^{\varepsilon}\bigl{(}i\psi,h_{ab}\gamma^{a}\varpi_{\mathfrak{m}}^{b}\wedge*D_{0}\psi\bigl{)}=\overbrace{S_{M}}^{\mathrm{Re}S_{D}}-\dfrac{1}{2}\int_{\mathscr{M}}h^{\varepsilon}\bigl{(}i\psi,h_{ab}\gamma^{a}\varpi_{\mathfrak{m}}^{b}\wedge*(\lambda\psi)\bigl{)}.$ (18)
3.2.2 The total action
The study of the full action $\tilde{S}_{T}=S_{G}+\tilde{S}_{M}$ is left for future work. In the following we will restrict ourselves to the study of the action $S_{T}=S^{\prime}_{G}+S_{M}=S_{G}+S_{M}-\int_{\mathscr{M}}\dfrac{vk}{2\pi^{2}\ell\ell^{\prime}}\eta_{ab}T^{a}\wedge\lambda\wedge\beta^{b}$.
@@ 3.2.2 The total action
$\dfrac{\delta\mathscr{L}^{\prime}_{G}[A,\beta,\lambda]}{\delta\beta^{c}}$ $=\dfrac{ek}{8\pi^{2}\ell\ell^{\prime}}(R^{ab}+\dfrac{k}{\ell\ell^{\prime}}\beta^{a}\wedge\beta^{b})\wedge\beta^{d}\varepsilon_{abcd}+\dfrac{k}{2\pi^{2}\ell\ell^{\prime}}(yR^{ab}\wedge\beta_{b}\eta_{ac}+2vT_{c}\wedge\lambda-v\beta_{c}\wedge d\lambda)$
@@
$\dfrac{\delta\mathscr{L}^{\prime}_{G}[A,\beta,\lambda]}{\delta A^{ab}}$ $=\dfrac{ek}{8\pi^{2}\ell\ell^{\prime}}T^{c}\wedge\beta^{d}\varepsilon_{abcd}+\dfrac{k}{2\pi^{2}\ell\ell^{\prime}}\bigl{(}yT_{a}-v\lambda\wedge\beta_{a}\bigr{)}\wedge\beta_{b}=\dfrac{1}{2}\mathfrak{s}_{ab}=-\dfrac{\delta\mathscr{L}_{M}[A,\beta,\lambda]}{\delta A^{ab}}$ (20)
$\dfrac{\delta\mathscr{L}^{\prime}_{G}[A,\beta,\lambda]}{\delta\lambda}$ $=-\dfrac{vk}{2\pi^{2}\ell\ell^{\prime}}T^{a}\wedge\beta_{a}=0=-\dfrac{\delta\mathscr{L}_{M}[A,\beta,\lambda]}{\delta\lambda}\ .$ (21)
Equation (21) implies either $v=0$ or $T^{a}\wedge\beta_{a}=0$.
For $v=0$, one then recovers the usual equations of motion associated to Einstein-Cartan gravity with a Holst action and bare cosmological constant term, and torsion $T$ linked to the spin density of matter.
In the other case, using the Hodge star operator lets us show that $T^{a}\wedge\beta_{a}=0$ is equivalent to requiring the axial vector part of torsion to vanish ($\tilde{t}^{a}=\dfrac{1}{6}T^{b}_{~{}\mu\nu}\varepsilon^{a~{}\mu\nu}_{~{}b}=0$) as is described in [28, Table 3]. The end result being that the new torsion $T^{a}_{~{}bc}=Z^{a}_{~{}bc}+\delta^{a}_{[b}t^{\vphantom{a}}_{c]}$ can be expressed solely in terms of $20$ independent components instead of the usual $24$. Here $Z^{a}_{~{}bc}$ is a $3$-indices tensor and $t_{c}$ is a $4$-dimensional vector with respectively $16$ and $4$ independent components.
@@
$\dfrac{ek}{8\pi^{2}\ell\ell^{\prime}}(R^{ab}+\dfrac{k}{\ell\ell^{\prime}}\beta^{a}\wedge\beta^{b})\wedge\beta^{d}\varepsilon_{abcd}+\dfrac{k}{2\pi^{2}\ell\ell^{\prime}}\bigl{(}yR^{ab}\wedge\beta_{b}\eta_{ac}-v\beta_{c}\wedge d\lambda\bigr{)}$ $=\tau_{c}$ (22)
$-\dfrac{vk}{\pi^{2}\ell\ell^{\prime}}\lambda\wedge\beta_{a}\wedge\beta_{b}$ $=\mathfrak{s}_{ab}$ (23)
@@
where equation (23) links spin density to the scalar gauge fields for dilation:
$\lambda_{c}=\lambda_{\mu}\beta^{-1\,\mu}_{~{}~{}\,\,\;\;\;c}=-\dfrac{\pi^{2}\ell\ell^{\prime}}{6vk}\mathfrak{s}^{ab}_{~{}~{}abc}\ .$ (25)
Using the relation $\gamma=\dfrac{2\alpha_{1}}{2\alpha_{2}-\alpha_{5}}$, given in [16, p13], we identify $\gamma=\dfrac{e}{r+2y}$, with $\gamma$ the Barbero-Immirzi parameter.
Provided we set $e=\dfrac{8\pi^{2}\alpha_{1}\ell\ell^{\prime}}{k}=\dfrac{3\pi}{2\Lambda_{0}G}$ and $\ell\ell^{\prime}=-\dfrac{3k}{\Lambda_{0}}$ where $\alpha_{1}=-\dfrac{1}{16\pi G}$, applying the Hodge star * to (22) and replacing $\lambda$ according to (25) gives us the Einstein equations modified by the Barbero-Immirzi part of the action with bare cosmological constant $\Lambda_{0}$ and with matter sources given by the energy-momentum tensor and a particular combination of derivatives of the spin density of matter (the latter resulting from the presence of the dilation 1-form $\lambda$ of the connection):
$G_{kc}+(\dfrac{1}{2\gamma}-\dfrac{\Lambda_{0}Gr}{3\pi})R^{ab}_{~{}~{}rs}\varepsilon^{rs}_{~{}~{}bk}\eta_{ac}+\Lambda_{0}\eta_{kc}=-\dfrac{1}{2\alpha_{1}}(\tau_{ck}+\dfrac{1}{12}\partial_{\mu}\mathfrak{s}^{ab}_{~{}~{}ab\nu}\varepsilon^{\mu\nu}_{~{}~{}\;ck}).$ (26)
Remark.
In this particular case the fact that Torsion is null implies that the Einstein tensor $G_{kc}$ is symmetric (see [29, p3] for more details). When taking into account that $\eta_{kc}$ is symmetric as well, the anti-symmetricity of $\partial_{\mu}\mathfrak{s}^{ab}_{~{}~{}ab\nu}\varepsilon^{\mu\nu}_{~{}~{}\;ck}$ implies that the energy-momentum tensor $\tau_{ck}$ sports an anti-symmetric part if $\partial_{\mu}\mathfrak{s}^{ab}_{~{}~{}ab\nu}\varepsilon^{\mu\nu}_{~{}~{}\;ck}\neq 0$. This confirms that we can consider a matter Lagrangian that depends on the spin connection $A$ as we explicitly did.
@@
$G_{kc}+(\dfrac{1}{2\gamma}-\dfrac{\Lambda_{0}Gr}{3\pi})R^{ab}_{~{}~{}rs}\varepsilon^{rs}_{~{}~{}bk}\eta_{ac}+\Lambda\eta_{kc}=-\dfrac{1}{2\alpha_{1}}(\tau_{M,ck}+\dfrac{1}{12}\partial_{\mu}\mathfrak{s}^{ab}_{~{}~{}ab\nu}\varepsilon^{\mu\nu}_{~{}~{}\;ck})$ (27)
with $\Lambda=\Lambda_{0}+\Lambda_{\text{vac}}=\overbrace{-\dfrac{3k}{\ell\ell^{\prime}}}^{\Lambda_{0}}\overbrace{-\dfrac{\rho_{\text{vac}}}{2\alpha_{1}}}^{\Lambda_{\text{vac}}}=-\dfrac{3k}{\ell\ell^{\prime}}-\dfrac{4\pi^{2}\ell\ell^{\prime}\rho_{\text{vac}}}{ek}=\Lambda_{0}+8\pi G\rho_{\text{vac}}$ the effective cosmological constant, and $\Lambda_{\text{vac}}$ the vacuum contribution to the cosmological constant.
@@
$G_{kc}$ $+(\dfrac{1}{2\gamma}-\dfrac{\Lambda_{0}Gr}{3\pi})R^{ab}_{~{}~{}rs}\varepsilon^{rs}_{~{}~{}bk}\eta_{ac}+\Lambda_{0}\eta_{kc}$
$=-\dfrac{1}{2\alpha_{1}}\Bigl{(}\tau_{ck}+\dfrac{1}{12}(\eta_{cs}\beta^{s}_{\nu}\partial_{\mu}+T_{c,\mu\nu})\bigl{(}\mathfrak{s}^{ab}_{~{}~{}abr}+(\dfrac{1}{2\pi G\gamma}-\dfrac{\Lambda_{0}r}{3\pi^{2}})T^{m}_{~{}rm}\bigr{)}\varepsilon^{\mu\nu r}_{~{}~{}~{}\,k}\Bigr{)}$ (28)
$\lambda_{r}$ $=\dfrac{1}{v}\bigl{(}\dfrac{\pi}{16\Lambda_{0}G_{0}}\overbrace{T^{c}_{~{}ab}\varepsilon^{ab}_{~{}~{}cr}}^{-6\tilde{t}_{r}=0}+\dfrac{y}{3}T^{c}_{~{}rc}+\dfrac{\pi^{2}}{2\Lambda_{0}}\mathfrak{s}^{ab}_{~{}~{}abr}\bigr{)}=\dfrac{1}{v}\bigl{(}\dfrac{y}{3}T^{c}_{~{}rc}+\dfrac{\pi^{2}}{2\Lambda_{0}}\mathfrak{s}^{ab}_{~{}~{}abr}\bigr{)}$ (29)